OpenStudy (anonymous):

the lengths of the side of a right triangle are x, x+3, and x+6. What is the length if the hypotenuse of this triangle

OpenStudy (anonymous):

x^2+(x+3)^2=(x+6)^2 x=9 hypotenuse is the longest so 9+6=15

OpenStudy (anonymous):

x x+3 x+6 x^2 (x+3)^2 (x+6)^2 c^2-a^2-b^2=0 1 4 7 1 16 49 32 2 5 8 4 25 64 35 3 6 9 9 36 81 36 4 7 10 16 49 100 35 5 8 11 25 64 121 32 6 9 12 36 81 144 27 7 10 13 49 100 169 20 8 11 14 64 121 196 11 9 12 15 81 144 225 0 10 13 16 100 169 256 -13 11 14 17 121 196 289 -28 12 15 18 144 225 324 -45 13 16 19 169 256 361 -64 14 17 20 196 289 400 -85 15 18 21 225 324 441 -108 16 19 22 256 361 484 -133 17 20 23 289 400 529 -160

OpenStudy (anonymous):

Plug your sides into the Pythagorean Theorem and then simplify and solve for x \[a^2+b^2=c^2\] Let a=x,b=x+3, and c=x+6 since x+6 has to be the longest side Therefore we will have \[x^2+x^2+6x+9=x^2+12x+36\] So now we simplify to get \[2x^2+6x+9=x^2+12x+36\] then move terms to set equation to 0 \[x^2-6x-27=0\] factor to get (x-9)(x+3)=0 x={-3,9} Since x is the length of one of the sides it cannot be negative so we can discard -3 as a solution, therefore we set our hypotenuse to be c=x+6 so c=9+6=15

OpenStudy (anonymous):

sashley0915 answer is correct given x+6 is the hyp